OpenStudy (yanasidlinskiy):

Please Help!!:D Rewrite the expression as a trigonometric function of a single angle measure.

OpenStudy (anonymous):

sec = 1/cos

OpenStudy (anonymous):

Is he correct @YanaSidlinskiy

OpenStudy (yanasidlinskiy):

Please stop. \[\frac{ \tan4\Theta +\tan2\Theta }{ 1-\tan 4\Theta \tan2\Theta } \]

OpenStudy (yanasidlinskiy):

Hate these things like crazy.

OpenStudy (anonymous):

u telling me to stop @YanaSidlinskiy

OpenStudy (yanasidlinskiy):

No, Maxwell:)

OpenStudy (yanasidlinskiy):

@paki

OpenStudy (paki):

oh... stay away form trig.... :(

OpenStudy (yanasidlinskiy):

I really need help. This is one of my questions on my assignment.

OpenStudy (paki):

@SolomonZelman please help here...

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \frac{\tan 4θ+\tan 2θ}{1-\tan 4θ\tan 2θ}\large{=\tan(4θ+2θ)=\tan(6θ)} }\)

OpenStudy (solomonzelman):

Like this, you want it, or everything in terms `tanθ` ?

OpenStudy (solomonzelman):

warning, in terms of `tanθ` will give a very awkward answer... so don't recommend .

OpenStudy (yanasidlinskiy):

That's what I was thinking. No. I don't need it in terms of tan. I've tried it and I know what it's like. I agree. It's definitely crazy:)

OpenStudy (yanasidlinskiy):

Thanx sooooooooooooooo much:)

OpenStudy (solomonzelman):

Anytime :) btw, my grade for these was 98 :)

OpenStudy (yanasidlinskiy):

OMG!!!!!!!!!!!!!!!!!!!!!!!!!!!!! I am soooooooo mad. Wanna teach me?

OpenStudy (solomonzelman):

What do you want to know?

OpenStudy (yanasidlinskiy):

How to solve these stupid things:)

OpenStudy (yanasidlinskiy):

Like examples.

OpenStudy (solomonzelman):

Well, I just remember the rules... I can only say that it is common sense after you know the identities. One tip for verifying trigonometric equations though, is to translate everything into sines and cosines (usually).

OpenStudy (yanasidlinskiy):

Ok, Well, then tell me the rules:) Hopefully I'll be smart like ya:)

OpenStudy (solomonzelman):

\(\huge\color{blue}{ \tan (a+b)=\frac{\tan(a)+\tan(b)}{1-\tan(a)\tan(b)} }\) \(\huge\color{blue}{ \tan (a-b)=\frac{\tan(a)-\tan(b)}{1+\tan(a)\tan(b)} }\)

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \sin (a±b)= \sin(a)\cos(b)± \sin(b)\cos(a) }\) \(\LARGE\color{blue}{ \cos (a±b)= \cos(a)\cos(b)∓ \sin(b)\sin(a) }\)

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \sin^2x+\cos^2x=1 }\)

OpenStudy (solomonzelman):

The rest can be derived from these.

OpenStudy (solomonzelman):

Well besides the half angle ones. You will always find the online.

OpenStudy (yanasidlinskiy):

I have a quick question. On the first rule like tan(a+b) and tan(a-b) will they always be in that form?

OpenStudy (solomonzelman):

Yes even if you have imaginary/complex angles.

OpenStudy (yanasidlinskiy):

Wow! Ok. I think I can write a whole book about you! explaining better than a book:) For real! Filled with informationxD!!!

OpenStudy (yanasidlinskiy):

Can we do an example of each?

OpenStudy (solomonzelman):

Anyone explains better than a book, books don't explain they are there to confuse you. Well, beside English and History books.

OpenStudy (solomonzelman):

besides

OpenStudy (yanasidlinskiy):

I agree. Definitely!:) *they are there to confuse*

OpenStudy (yanasidlinskiy):

And..Can we do an example of each if you don't mind?

OpenStudy (solomonzelman):

Okay, but you will need to do it. And the questions won't be telling which rule to use. Ready ?

OpenStudy (yanasidlinskiy):

Umm.. Ok. I guess so. You have to check it though too. Ok?

OpenStudy (solomonzelman):

Verify the rule below. \(\normalsize\color{blue}{ \cos(2θ)=\cos^2θ - \sin^2θ }\)

OpenStudy (yanasidlinskiy):

Ok. Is it this? \(\normalsize\color{blue}{ \cos (a±b)= \cos(a)\cos(b)∓ \sin(b)\sin(a) }\)

OpenStudy (solomonzelman):

yes, so the rule is \(\normalsize\color{blue}{ \cos(a+b)=\cos(a) \cos(b) - \sin(b) \sin(a) }\)

OpenStudy (solomonzelman):

use it:)

OpenStudy (yanasidlinskiy):

Ok. Give me a sec. See what I can get:)

OpenStudy (solomonzelman):

sure

OpenStudy (yanasidlinskiy):

I dunno:( I don't know where to plug them in.

OpenStudy (yanasidlinskiy):

Wait..Wait.. I *think*

OpenStudy (solomonzelman):

Hint: `2θ = θ + θ `

OpenStudy (yanasidlinskiy):

I'm just curious..Is the answer going to be 0?

OpenStudy (solomonzelman):

You just need to show that `cos(2θ)= cos²θ - sin²θ`

OpenStudy (yanasidlinskiy):

Umm. Ok. well.. cos\[\cos(0\pm0) = \cos(0)^{2}\cos(0)^{2}\sin(2)^{2}\sin(0)\] I really don't know:) lol. At least I gave it a try:) Show me:)

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \cos(2θ)=\cos(θ+θ)=\cos(θ)\cos(θ)- \sin(θ)\sin(θ)= }\) \(\normalsize\color{blue}{ \cos^2θ- \sin^2θ. }\)

OpenStudy (solomonzelman):

Makes sense ?

OpenStudy (yanasidlinskiy):

Yes. Better than mine!!!!:D Wow!! Some more? Are you tired?

OpenStudy (solomonzelman):

I can give you one more question. Show that \(\normalsize\color{blue}{ \cos~90 = 0. }\)

OpenStudy (solomonzelman):

Post each step as you go,

OpenStudy (yanasidlinskiy):

Will I use this formula? \(\normalsize{ \sin^2x+\cos^2x=1 }\)

OpenStudy (solomonzelman):

No

OpenStudy (solomonzelman):

90 = 45 + 45 (hint)

OpenStudy (yanasidlinskiy):

Will it be the same formula as we used for the other question?

OpenStudy (solomonzelman):

Yes.... and you will agree that sin(45)=soc(45) right ? Now, do the proof

OpenStudy (solomonzelman):

sin(45)=cos(45) I meant

OpenStudy (yanasidlinskiy):

Ok. let me try.

OpenStudy (solomonzelman):

sure, but on your way, do not forget to prove that sin(45)=cos(45), and find their values, and to plug them in.

OpenStudy (yanasidlinskiy):

That's a bummer. For real.

OpenStudy (yanasidlinskiy):

Do i need to show that cos 90=0 first?

OpenStudy (solomonzelman):

showing that cos(90) will not be possible before you show that cos(45)=sin(45)

OpenStudy (solomonzelman):

showing that cos(90)=0 will not be... (I meant)

OpenStudy (yanasidlinskiy):

Ok. Here's what I got: \[\cos(90)=\cos(45+45)=\cos(45)\cos(45)-\sin(45)\sin(45)=\cos ^{45} -\sin ^{45}\] Am I wrong? Ughh. You're going too hard one meh.xD

OpenStudy (yanasidlinskiy):

I think I messed up, for real. And probably BIG time too.

OpenStudy (solomonzelman):

Well, it is equal to `cos²(45) - sin²(45)`

OpenStudy (solomonzelman):

Now, show that sin(45) = cos(45) Draw a 45-45-90 and the proportions of it

OpenStudy (yanasidlinskiy):

So it was only there \(\huge{\uparrow}\) that I messed up?

OpenStudy (solomonzelman):

yes, you wrote 45's as powers, but 45s are angles and they are to the second power each.

OpenStudy (yanasidlinskiy):

Whew! I seriously am startin to understand it now:) And I'm not sure how I would do that.

OpenStudy (solomonzelman):

Draw a 45-45-90 triangle, and write the length of each side in terms of S.

OpenStudy (yanasidlinskiy):

Does it matter where I put the numbers?

OpenStudy (yanasidlinskiy):

And what do you mean by "terms of S?" Like Sin?

OpenStudy (solomonzelman):

Put the numbers next to the sides, because they are to lavel the side lengths. No, "S" is just a variable, can be "x", "a" "b" or any other variable.

OpenStudy (solomonzelman):

They are to LABEL the side lengths (sorry for the typo)

OpenStudy (yanasidlinskiy):

??